Most lotteries are order-insensitive, and typically present the results in ascending numeric order. The actual draw often happens in some order (e.g. numbered balls being drawn from a hopper), and it'd be even more unusual if the numbers were actually drawn in consecutive order, but drawing 5-2-3-1-4 would typically be presented as 1-2-3-4-5 and would still be remarked upon as unusual.
Doing this, your final figures should be 12.1% and 96.8%.
For the expected number of times it would happen though, you are right, it would be three or four times per year on average for the second case.
Edit: What I did was take all the possible combinations of 5 balls (5!) by the number of different sets that could be drawn (50-5, based on lowest number), over the total possible draws (50!). I think perhaps what that does it not account for overlap between sets (1-5 and 2-6), inflating the number somewhat, which is why it's a bit more than twice the probability than you got for any possible sequence of 5.
function rn() {
return parseInt(Math.random() * 100);
}
function drawing(){
let result = [rn(), rn(), rn(), rn(), rn()];
return result.sort();
};
function sequential(arr){
for(let i = 0; i < arr.length-1; i++){
if(arr[i] + 1 != arr[i+1]){
return false
}
}
return true
}
let counter = 0;
for (let i = 0; i < 10000000; i ++){
if (sequential(drawing())){
counter ++;
}
};
console.log(counter);
16
Mathematically, it seems like you'd need to draw one of five numbers from the range, then one of four numbers, then one of three... so the likelihood would be 5/100 * 4/100 * 3/100 * 2/100 * 1/100 = 0.000000012. Although those odds don't seem to line up with the javascript I posted.
Edit: but if the numbers don't have to be drawn in order (e.g. 8-5-9-6-7 is OK), then the odds are much higher still:
(With the initial "2" term accounting for 4 consecutive numbers on either side of the initial pick -- though I'm not sure I've got that entirely right?) Then it would happen three to four times a year. Even with a 6th ball drawn separately out of 20, that's still a 17% chance happening somewhere in the world in a year, given 1,000 daily draws.